Horace the duck is taking a bath.  He fills a tub with 45 L of water at 46 ℃. This is a little too warm for Horace, so he thinks to add some water at 10℃, to bring the temperature of the bath down to 35℃.  How much water will Horace have to add to the tub?  If the tub will hold 55 L of water will he be able to fill the bath without it overflowing?


1
Expert's answer
2021-06-07T17:19:41-0400
"\rho V'c(T-T_1)=\rho Vc(T_2-T)\\(35-10)V'=45(46-35)\\V'=20\ L"

"V+V'=45+20=65\ L"

The tub will not be able to fill the bath without it overflowing.